Matrix Inc. calculates cost for an equivalent unit of production using the FIFO method. Data for July: Work-in-process inventMultiple Choice $3.10. $3.00. $2.00. $2.40. $3.20.
Answer #1
Conversion costs
To complete Beginning Work in Process inventory 18000 =36000*(1-50%)
Units started and completed 66000 =102000-36000
Ending Work in Process inventory 12000 =24000*50%
Equivalent units 96000
Manufacturing costs for July 288000
/ Equivalent units 96000
Cost per Equivalent unit 3.00
$3.00 is correct option